OpenStudy (anonymous):

Solve for the variable in this equation. x/2+4/5=9/12

Parth (parthkohli):

30x/60 + 48/60 = 45/60 30x/60 = -3/60 30x = -3 x = -3/30 x = -1/10

Parth (parthkohli):

Thanks for fanning :)

OpenStudy (anonymous):

NP:)